OpenStudy (itsmichelle29):

What is the equation of a horizontal line passing through the point (2, 10)? y = 12 x = −8 y = 10 x = 2

OpenStudy (itsmichelle29):

is it x=2 ?? im not sure

OpenStudy (lυἶცἶ0210):

Horizontal line means it's (0,0)

OpenStudy (itsmichelle29):

so its x=2 ?? yes

OpenStudy (lυἶცἶ0210):

So just plug in (0, 0) to get a slope: \(\LARGE \frac{ 10-0}{2-0} \) gives use what? And nope

OpenStudy (lυἶცἶ0210):

Whoops, mess up there, one sec

OpenStudy (itsmichelle29):

5

OpenStudy (itsmichelle29):

????

OpenStudy (imtiaz7):

to look at it is that a horizontal line has a slope of zero, so using the point-slope form of the straight line

OpenStudy (imtiaz7):

and in horizontal the slope is always zero do you got it @itsmichelle29

OpenStudy (imtiaz7):

so y=10

OpenStudy (itsmichelle29):

and can it not be x=2

OpenStudy (imtiaz7):

because as we apply here point slope form y-y1=m(x-x1) y-10=0(x-2) y-10=0 and then y=10

OpenStudy (lυἶცἶ0210):

Yes, the slope is zero, but the coordinates are not always (0, 0), that's where I messed up, sorry. Since it's (2, 10) and it's passing through it, being horizontal it'll be ( 0, 10)

OpenStudy (itsmichelle29):

ohh lol okay i get it now thanks

OpenStudy (anonymous):

its y=10

OpenStudy (imtiaz7):

this question is good @itsmichelle29 and interesting one